OpenStudy (anonymous):

if dy/dx = 2xy^2, and y(-1) = 2, find y(2)

OpenStudy (raffle_snaffle):

separable differential equation

OpenStudy (raffle_snaffle):

put x on one side and y on the other side of the equals sign then integrate.

OpenStudy (anonymous):

I got as far as \[2\int\limits(y^2)dy=\int\limits(1/x)dx\] then I integrated to get \[(2/3)y^3 = \ln \left| x \right|+ C\] what now?

OpenStudy (raffle_snaffle):

Since you are given y(-1) = 2 you can solve for C

OpenStudy (raffle_snaffle):

X = -1 and y = 2

OpenStudy (anonymous):

I solved for C and got (16/3) but my problem was in later plugging in y(2), the answer had to be either (a)4e^3 (b) -4e^3 (c) -3/2 (d) -2/5 or (e) -1/4

OpenStudy (raffle_snaffle):

well we can' t take ln(-1)

OpenStudy (raffle_snaffle):

ln(-1) = 3.14159265 i

OpenStudy (raffle_snaffle):

@robtobey please help OP.

OpenStudy (loser66):

Your calculation is not correct

OpenStudy (raffle_snaffle):